OpenStudy (anonymous):

Fan and medal! Find the solution to the system of equations by using either graphing or substitution. y = 2x - 1 and y = x + 3

OpenStudy (gorv):

by substitution

OpenStudy (gorv):

y=x+3 here value of y is =x+3 put this value in place of y in the second equation

OpenStudy (gorv):

y=2x-1 replace y by x+3 here

OpenStudy (gorv):

@doglover789